# The 4D Human Embryonic Brain Atlas: spatiotemporal atlas generation for rapid anatomical changes using first-trimester ultrasound from the Rotterdam Periconceptional Cohort

### v1

@@ 5.4 Experiment 4: Influence of maternal BMI on morphology
Applying the $5\%$ FDR correction, we found that $p<0.027$ was significant. Using $\delta=3$ from Experiment 1, the GA window became 7 days. Figure 12 shows significant voxels in the mid-sagittal and mid-axial plane for each gestational week after performing VBM in red. The atlas and delineated regions correspond to the middle day of each week (days 59, 66, 73, 80, 87). Note that most significant differences occurred within or near the delineated regions. Starting at week 9, these significant differences were present in the cerebellum in the mid-axial plane, which was expected since in previous work a negative association was found between maternal BMI and trans-cerebellar diameter (Koning et al., 2016). There were also significant differences in non-delineated regions. For example in weeks 8, 11, and 12 in the axial plane, there was a hyper-intense region with significant p-values close to the fourth ventricle near the developing brain stem.
@@ 6 Discussion
Our method introduced a time-dependent initial atlas and penalized deviations from it, preserving age-specific anatomy during rapid development. Building on Dalca et al. (2019), we made two key contributions: using a time-dependent initial atlas instead of a single one for all gestational days, and constraining the atlas generation by penalizing deviations. In the ablation study, we found that both contributions were essential for an anatomically accurate atlas, while omitting them resulted in an atlas with incorrect anatomy.
@@
We evaluated the resulting 4D Human Embryonic Brain Atlas in terms of volumetric growth and anatomy. We found that the HV-curve of the atlas had an average relative difference of $12.2\%$ with respect to the curve obtained in previous research by Koning et al. (2016). This difference may partially be explained by the study population used to create the HV-curve. Namely, IVF/ICSI pregnancies were included, which has been shown to influence volumetric growth (Koning et al., 2016). In terms of anatomy, the visual comparison between the 4D Human Embryonic Brain Atlas and the 3D Embryo Atlas (de Bakker et al., 2016) demonstrated a similar appearance of brain structures and ventricles. The comparison to the Fetal Brain Atlas (Namburete et al., 2023) showed the same brain structures and ventricle shaped in subject-specific ultrasound images and the atlases, indicating that the 4D Human Embryonic Brain Atlas reflects the structural information present in the subject-specific ultrasound images similar to the Fetal Brain Atlas. Finally, using VBM, we assessed whether the atlas could capture known differences in morphology. To this end, we studied the influence of maternal overweight and obesity during early pregnancy on first trimester brain development. Using VBM, we found that the cerebellum signiticantly differs in morphology in case of a high maternal BMI, which is in line with previous research (Koning et al., 2017).
In future work, the VBM study could be improved since we performed a statistical test in every voxel (> 2 million) within the brain. This may lead to false positives despite the correction for multiple testing, and make the results hard to interpret, especially in non-delineated regions. A possible direction for future research is to use dimensionality reduction techniques, such as PCA, to extract the relevant morphological variations among subjects (Kurita, 2019). Dimensionality reduction techniques reduce the risks of false positives and may aid the interpretation of the results. Another limitation of the VBM study is that we did not model the longitudinal relation between images acquired at multiple points during the first trimester.
@@
When comparing the 4D Human Embryonic Brain Atlas to the Fetal Brain Atlas (Namburete et al., 2023), we found that the Fetal Brain Atlas showed more details and clearer structures. This might be explained by the several additional image processing steps taken, including speckle noise reduction and manually annotated landmarks, to improve the sharpness of brain structures in the Fetal Brain Atlas. In future work, besides including additional (manual) image processing steps, generative approaches such as latent diffusion models can be explored to improve atlas generation since they often produce sharper and more detailed images than registration-based approaches (Rombach et al., 2022).
A potential direction for future research to enhance visible details and brain structures is to refine the creation of the initial atlas. In the current approach, all images within the $\delta$-window are treated as equally important when generating the initial atlas. A possible extension could involve using (weighted) kernel regression, a technique commonly applied in non-learning-based methods (Kuklisova-Murgasova et al., 2011; Serag et al., 2012; Gholipour et al., 2007).
@@
Regarding the clinical impact of the 4D Human Embryonic Brain Atlas, firstly, as demonstrated, the atlas can be used to analyze the influence of paternal and maternal conditions and lifestyle behaviors on brain development. Using a similar analysis, early detection of (congenital) anomalies and adverse growth patterns may be performed by comparing new subjects to the atlas. Secondly, the atlas holds significant value as a visualization aid for soon-to-be parents in clinical practice. It offers a unique and detailed perspective on the progression of early brain growth. Research has shown that confirmation of healthy development as shown during a standard ultrasound scan promotes parental attachment (Walsh et al., 2014; Boukydis et al., 2009; Lapaire et al., 2007; O’Malley et al., 2020). Moreover, studies investigating novel visualization techniques of first trimester ultrasonography, such as offering an additional 3D VR visualization or using real-time ultrasound, observe a positive parental response and lead to reduced anxiety (Pietersma et al., 2022; Lapaire et al., 2007; Popovici et al., 2019).
In conclusion, the proposed deep learning approach for atlas generation can accurately capture age-specific morphological features in the case of rapid brain development during the first trimester. Using the proposed model, we generated the 4D Human Embryonic Brain Atlas, the first spatiotemporal atlas of the human embryonic and fetal brain, covering days 56 to 90 of pregnancy. The 4D Human Embryonic Brain Atlas provides a unique, detailed, and accurate picture of brain development during the first trimester. This ultimately may lead to earlier and better detection, prevention, and treatment of neurodevelopmental disorders.

### v2

@@ 5.4 Experiment 4: Influence of maternal BMI on morphology
Applying the $5\%$ FDR correction, we found that $p<0.027$ was significant in all the analyzes. Using $\delta=3$ from Experiment 1, the GA window became 7 days. Figure 12 shows significant voxels in the mid-sagittal and mid-axial plane for each gestational week after performing VBM in red when comparing normal maternal BMI to high maternal BMI. The atlas and delineated regions correspond to the middle day of each week (days 59, 66, 73, 80, 87). Note that most significant differences occurred within or near the delineated regions. Starting at week 9, these significant differences were present in the cerebellum in the mid-axial plane, which was expected since in previous work a negative association was found between maternal BMI and trans-cerebellar diameter (Koning et al., 2016). There were also significant differences in non-delineated regions. For example in weeks 8, 11, and 12 in the axial plane, there was a hyper-intense region with significant p-values close to the fourth ventricle near the developing brain stem.
@@
We additionally performed subgroup analyses comparing normal BMI vs. overweight, normal BMI versus obese, and overweight versus obese. The normal versus overweight results largely overlapped with the combined high-BMI analysis and also showed additional differences near the fourth ventricle at 12 weeks (Supplementary Figures 15A and B and 16A and B). The normal versus obese comparison revealed differences in the mesencephalon that were less prominent when overweight and obesity were combined (Supplementary Figures 15A and C and 16A and C). The overweight versus obese analysis most clearly highlighted subgroup-specific effects, showing differences in the mesencephalon at week 10 and the fourth ventricle at week 12, consistent with the separate analyses (Supplementary Figures 15D and 16D).
@@ 6 Discussion
Our method introduces a time-dependent initial atlas and penalizes deviations to preserve age-specific anatomy during rapid first trimester development. While building on Dalca et al. (2019), which was evaluated in the context of Alzheimer’s disease and characterized by relatively subtle anatomical changes, our approach explicitly models the pronounced and rapid anatomical variations of early brain development. The ablation study demonstrated that both adaptations were essential for anatomical accuracy, with their omission resulting in incorrect atlas anatomy, highlighting the necessity of these contributions for generating age–specific atlases that maintain anatomical consistency over time.
@@
We evaluated the resulting 4D Human Embryonic Brain Atlas in terms of volumetric growth and anatomy. We found that the HV-curve of the atlas had an average relative difference of $12.2\%$ with respect to the curve obtained in previous research by Koning et al. (2016). This difference may partially be explained by the study population used to create the HV-curve. Namely, IVF/ICSI pregnancies were included, which has been shown to influence volumetric growth (Koning et al., 2016). In terms of anatomy, the visual comparison between the 4D Human Embryonic Brain Atlas and the 3D Embryo Atlas (de Bakker et al., 2016) demonstrated a similar appearance of brain structures and ventricles. The comparison to the Fetal Brain Atlas (Namburete et al., 2023) showed the same brain structures and ventricle shaped in subject-specific ultrasound images and the atlases, indicating that the 4D Human Embryonic Brain Atlas reflects the structural information present in the subject-specific ultrasound images similar to the Fetal Brain Atlas. Finally, using VBM, we assessed whether the atlas could capture known differences in morphology. To this end, we studied the influence of maternal overweight and obesity during early pregnancy on first trimester brain development. Using VBM, we found that the cerebellum significantly differs in morphology in case of a high maternal BMI, which is in line with previous research (Koning et al., 2017).
Several limitations should be considered. In the VBM study, statistical tests were performed at every voxel (>2 million) within the brain. This may lead to false positives despite correction for multiple testing and make the results difficult to interpret, particularly in non-delineated regions. Furthermore, the analysis was performed on the entire embryonic head rather than the isolated brain, following the definition of the HV used in previous work (Koning et al., 2016). As non-brain tissue cannot yet be clearly separated during the first trimester and no standardized segmentation method exists, this ensured anatomical completeness but may have reduced interpretability in some regions. Future work should aim to segment individual brain regions to enable more localized analyses. In addition, the application of dimensionality reduction techniques, such as PCA (Kurita, 2019), could reduce the risk of false positives and facilitate interpretation by extracting the most relevant morphological variations among subjects.
A further consideration is the grouping of overweight and obese participants in the primary analysis. Our subgroup analyses showed that overweight and obesity separately led to partly distinct spatial patterns, some of which were less apparent when the categories are combined. These differences should be interpreted with caution given the limited sample size, particularly in the obesity group, and future studies are needed to validate and further explore these differences.
Another limitation is that we did not model the longitudinal relationship between images acquired at multiple time points during the first trimester. Although several participants underwent repeated scans, these were treated as independent samples. Future work could incorporate longitudinal information directly into the atlas generation process to better capture individual growth trajectories. Existing atlas methods listed in 1 that include longitudinal data still treat these images as cross-sectional, underscoring the challenges of capturing individual growth trajectories. Approaches that account for intra-subject trajectories often use a two-stage process: first constructing subject-specific atlases by registering intra-subject data, then creating a population-level atlas. This typically requires carefully designed datasets with imaging data available for all subjects at each time point (Liao et al., 2012; Li et al., 2015). A promising direction for future work is to combine the proposed time-dependent initial atlas with an implicit neural representation framework, enabling direct modeling of subject-specific developmental trajectories (Chen et al., 2025).
@@
Compared with the Fetal Brain Atlas (Namburete et al., 2023), our atlas shows less detail and sharper structures. This might be explained by the several additional image processing steps taken, including speckle noise reduction and manually annotated landmarks, to improve the sharpness of brain structures in the Fetal Brain Atlas. In future work, besides including additional (manual) image processing steps, generative approaches such as latent diffusion models can be explored to improve atlas generation since they often produce sharper and more detailed images than registration-based approaches (Rombach et al., 2022). Another potential direction for future research to enhance visible details and brain structures is to refine the creation of the initial atlas. In the current approach, all images within the $\delta$-window are treated as equally important when generating the initial atlas. A possible extension could involve using (weighted) kernel regression, a technique commonly applied in non-learning-based methods (Kuklisova-Murgasova et al., 2011; Serag et al., 2012; Gholipour et al., 2007).
@@
Finally, regarding the clinical impact of the 4D Human Embryonic Brain Atlas, firstly, as demonstrated, the atlas can be used to analyze the influence of paternal and maternal conditions and lifestyle behaviors on brain development. Using a similar analysis, early detection of (congenital) anomalies and adverse growth patterns may be performed by comparing new subjects to the atlas. Secondly, the atlas holds significant value as a visualization aid for soon-to-be parents in clinical practice. It offers a unique and detailed perspective on the progression of early brain growth. Research has shown that confirmation of healthy development as shown during a standard ultrasound scan promotes parental attachment (Walsh et al., 2014; Boukydis et al., 2009; Lapaire et al., 2007; O’Malley et al., 2020). Moreover, studies investigating novel visualization techniques of first trimester ultrasonography, such as offering an additional 3D VR visualization or using real-time ultrasound, observe a positive parental response and lead to reduced anxiety (Pietersma et al., 2022; Lapaire et al., 2007; Popovici et al., 2019).
7 Conclusion
We proposed a deep learning approach for spatiotemporal atlas generation capable of accurately capturing age-specific morphological features during rapid first trimester brain development. Using this approach, we generated the 4D Human Embryonic Brain Atlas, the first spatiotemporal atlas of the human embryonic brain covering days 56 to 90 of gestation. The atlas provides a framework for quantitative analysis of early brain development, enabling the study of parental and lifestyle factors, supporting early detection of congenital anomalies and abnormal growth patterns, and serving as a clinically valuable visualization aid for parents and healthcare providers. Overall, the 4D Human Embryonic Brain Atlas offers a detailed and accurate picture of early brain development, which may contribute to earlier detection, prevention, and treatment of neurodevelopmental disorders while facilitating further research in this critical developmental period.
@@ Acknowledgements
Appendix A Supplementary material 1: Influence of $\delta$ during gestational week 8
Figure 14: Atlas $A_{t}$ in gestational week 8 for different values of $\delta$. For $\delta\geq 4$ the choroid plexus appears starting at day 56, which is anatomically incorrect. All images are shown in the mid-sagittal plane. The arrows indicate relevant visible brain structures. LV = lateral ventricle, 3V = third ventricle (cavity of the diencephalon), M = cavity of the mesencephalon, 4V = fourth ventricle (cavity of the rhombencephalon), CP = choroid plexus, CER = cerebellum.
Appendix B Supplementary material 2: Subgroup analysis experiment 4
Figure 15: Percentage of significantly different voxels within the delineated ventricles and brain structures at different gestational weeks for the VBM subgroup analysis in Experiment 4. LV = lateral ventricle, 3V = third ventricle (cavity of the diencephalon), M = cavity of the mesencephalon, 4V = fourth ventricle (cavity of the rhombencephalon), CP = choroid plexus, CER = cerebellum.
Figure 16: Visualization of the comparison between brain morphology for every subgroup analysis using VBM in Experiment 4. The red voxels indicate the p-value for significant structural volumetric differences (p<0.027) and the black arrows highlight the prominent differences between the subgroups. The atlas of the middle of every gestational week is shown. LV = lateral ventricle, 3V = third ventricle (cavity of the diencephalon), M = cavity of the mesencephalon, 4V = fourth ventricle (cavity of the rhombencephalon), CP = choroid plexus, CER = cerebellum.